# Decentralized and Uncoordinated Learning of Stable Matchings: A Game-Theoretic Approach

### v1

@@ V Global Learning of Stable Matchings in General Markets
A pure strategy profile $x$ is a good state if all the men matched under $x$ are at their best responses. [^8] In a pure strategy profile $x=(x_{m},m\in M)$, a matched man is the one whose pure strategy $x_{m}$ has exactly one coordinate equal to 1 (say $w$th coordinate), and all other coordinates are zero, and no other man of higher preference proposes to $w$, i.e., $x_{kw}=0\ \forall k>_{w}m.$ In other words, a good state $x$ refers to a partial matching in which the matched men are at their best responses.
@@
According to the payoff structures, the best response of player $m$ at time $t$ is to propose to his most preferred achievable woman $w^{\prime}$, i.e., among all women who rank $m$ higher or equal to their match under $x^{t-1}$, $m$ proposes to the woman $w^{\prime}$ which brings him the highest utility $\mu_{mw^{\prime}}$. Note that in particular $\mu_{mw^{\prime}}>\mu_{mw}$. Thus, the first time $\tau\leq t-1$ when man $m$ proposed and got matched to woman $w$, the woman $w^{\prime}$ was his more preferred choice and the reason why $m$ did not propose to $w^{\prime}$ at time $\tau$ was because $w^{\prime}$ was not achievable, i.e., $x^{\tau-1}_{m^{\prime}w^{\prime}}=1$ for some man $m^{\prime}>_{w^{\prime}}m$. Since $w^{\prime}$ was not achievable for $m$ in $x^{\tau-1}$ but she is achievable in $x^{t-1}$ implies that $r_{w^{\prime}}(x^{t-1})\!<\!r_{w^{\prime}}(x^{\tau-1})$. Therefore, $r_{w^{\prime}}(x^{\tau^{\prime}})\!<\!r_{w^{\prime}}(x^{\tau^{\prime}-1})$ for some $\tau\leq\tau^{\prime}\leq t-1$, contradicting the choice of $t$.  Q.E.D.
In fact, one can view the DA algorithm of [1] as a special version of the above lemma in which all the players are selected to best respond at each round, and the initial good state is the empty matching. Lemma 6 shows that the stable matching game admits an ordinal potential function if the game starts from a good state (we refer to Appendix A-F for another interpretation of this potential function). However, that does not imply that the stable matching game is an ordinal potential game [21], which requires the existence of a potential function whose changes are aligned with players’ payoff changes from any initial state. Nonetheless, the existence of such a “weak” potential function allows us to show the weak acyclicity of the stable matching game as defined next.
@@
Using the above definitions, we are now ready to state the main result of this section, which provides a decentralized and uncoordinated algorithm for learning stable matchings in general markets with unknown preferences. The algorithm is an adaptation of the sample experimentation algorithm for weakly acyclic games [23] to the stable matching game. Intuitively, each player $m$ goes through a sequence of phase-dependent exploration/exploitation. At the end of each episode, player $m$ evaluates his payoff for each action he has taken during the last episode. He then updates his baseline action (an action that brings him the highest reward in the past episode up to some tolerance) and sticks to that baseline action most of the time in the next episode while still exploring new actions with a small probability. The detailed description of the algorithm is summarized in Algorithm 2.
@@
  • During the $s$th episode, player $m$ selects his baseline action obtained at the end of the previous episode with probability $1-\epsilon$ or explores a new uniformly sampled action with probability $\epsilon$, i.e.,
@@
  • At the end of episode $s$, player $m$ evaluates his average utility when playing each of his actions as
@@
Next, we note that any pure NE in the stable matching game is a strict NE. [^9] In a weakly acyclic game, an arbitrarily high probability convergence can be guaranteed to a strict NE [23]. If we do not have strictness property, the dynamics may oscillate between two nonstrict NE points with a positive probability. This follows directly by the fact that there are no ties between men’s preferences in the stable matching game. More precisely, let $\alpha^{*}$ be a pure NE (stable matching) and assume $u_{m}(\alpha^{*})=\mu_{mw}$. For any $\alpha_{m}\neq\alpha^{*}_{m}$, either $u_{m}(\alpha_{m},\alpha^{*}_{-m})$ equals 00 or $\mu_{mw^{\prime}}$ for some $w^{\prime}\neq w$. Because $\alpha^{*}_{m}$ is the best response of player $m$ at the NE and there are no ties between man $m$’s preferences, we must have $\mu_{mw^{\prime}}<\mu_{mw}$. Since $0<\mu_{mw}$, in either case we have $u_{m}(\alpha_{m},\alpha^{*}_{-m})<u_{m}(\alpha^{*}_{m},\alpha^{*}_{-m})$.
@@
VI Matching Markets with Stronger Information Feedback
So far, we have assumed that men receive the minimum amount of information feedback as they interact with the market; namely, men only get to observe whether their proposal gets rejected or, if matched, they observe a random realization of their preference. In this final section, we consider the learning problem from a new perspective by addressing the following question: Is there stronger information feedback that will allow us to derive the market faster globally to a stable matching in a decentralized and uncoordinated fashion? Therefore, we address the problem of learning a stable matching from an information feedback design perspective by showing that if the players are allowed to receive a bit more information as feedback, then the stable matching game will become a monotone game, hence allowing a wide range of decentralized and uncoordinated learning algorithms for obtaining its NE points.
@@ VI Matching Markets with Stronger Information Feedback
Consider the regularized stable matching game with list feedback, where each player aims to maximize its regularized payoff function given by $U_{m}(x_{m},x_{-m}):=u_{m}(x_{m},x_{-m})-\frac{\beta}{2}\|x_{m}\|^{2}$. Then, for any two strategy profiles $x^{\prime},x$, we have
@@
Remark 6
The monotonicity property in Theorem 8 holds when the regularizer parameter $\beta$ is sufficiently large. The regularizer implies that players tend to maximize their approximate payoff functions. Thus, instead of learning a NE for the payoff functions (68), they learn an $\epsilon$-NE where the approximation error $\epsilon$ grows as $\beta$ increases. As a result, applying decentralized learning algorithms to learn a NE of the regularized game only guarantees an $\epsilon$-stable matching in the sense that the players will achieve a solution in which each player would be $\epsilon$ worse off than what he could achieve in a stable matching. Therefore, the regularizer parameter trades off between the market’s convergence speed and the accuracy of the learned stable matching.
Finally, we note that without imposing any assumption on the agents’ preferences, one cannot expect the stable matching game to be a monotone game. The reason is that monotone games imply the global uniqueness of a NE [24]. However, by Proposition 7, any stable matching is a pure NE, and for general matching markets, there could be many stable matchings under arbitrary preferences. Moreover, the assumption of waiting list feedback is not necessary for faster convergence in any general market; it is merely one choice that we consider in this work, and there could be other types of information feedback that allow faster convergence. Nevertheless, our game-theoretic approach to the stable matching problem provides valuable insights on how to properly design the utilities and the information feedback to devise faster decentralized and uncoordinated algorithms for learning stable matchings.
VII Conclusions
In this work, we considered the problem of learning stable matchings with unknown preferences in a fully decentralized and uncoordinated manner. Using a game-theoretic formulation, we showed how to leverage the rich literature for learning NE in noncooperative games to derive the matching market to its stable points in a principled way. In particular, we established several global and local convergence results with performance guarantees for learning stable matchings in hierarchical markets, general markets, and under weak/strong information feedback. Our results provide new insights for learning stable matchings through NE learning in noncooperative games and bridge the discrete problem of learning a stable matching with learning NE in continuous-action games.
This work opens several avenues for future research. One interesting direction is to characterize an information-theoretic limit below and above which efficient learning of a stable matching in a fully decentralized and uncoordinated fashion may or may not be possible. Moreover, our waiting list feedback is merely an initial attempt to design faster decentralized and uncoordinated algorithms under richer information feedback. Extending our results to include other types of information feedback and studying their effect on the convergence speed of the market to its stable points would be another interesting research direction.
Appendix A Supplementary materials
A-A Freedman’s Concentration Inequality for Martingales
@@
The proof can be found in [30, Theorem 1.6].  Q.E.D.
A-B Proof of Theorem 2
Proof:
$\prod_{k>_{w_{1}}m}(1-x_{kw_{1}})=\prod_{k>_{w_{2}}m}(1-x_{kw_{2}})=1,$
which together with the above argument implies that
A-C Proof of Lemma 4
@@ A-D Proof of Lemma 5
A-E Proof of Theorem 5
Proof:
For any $w^{\prime}\neq w$, by taking limit from the above relation as $t\to\infty$, we get $\limsup_{t\to\infty}\eta^{t}A^{t-1}\leq\ln(\frac{X^{*}_{mw}}{X^{*}_{mw}})+3\leq O(1)$. [^11] Here, if $X^{*}_{mw^{\prime}}=0$, by convention we define $\ln(\frac{X^{*}_{mw^{\prime}}}{X^{*}_{mw}})=-\infty$. Now let $\kappa:=v_{mw^{\prime}}(X^{*})-v_{mw}(X^{*})$. We claim $\kappa\leq 0$. Otherwise, if $\kappa>0$, then for any sufficiently large $t>\bar{t}$ we must have $A^{t-1}>t\kappa/2$. Given the choice of stepsize, we obtain $\lim_{t\to\infty}\eta^{t}A^{t-1}=\lim_{t\to\infty}t\eta^{t}\kappa/2=\infty$, which contradicts the fact that $\limsup_{t\to\infty}\eta^{t}A^{t-1}\leq O(1)$. Thus, $\kappa\leq 0$, which shows that if $X^{*}_{mw}>0$, then $v_{mw}(X^{*})=\max_{w^{\prime}}v_{mw^{\prime}}(X^{*})$, i.e., $X^{*}$ is a NE.  Q.E.D.
A-F An Alternative Form for the Potential Function in Lemma 6
One can represent the potential function in Lemma 6 directly for womens’ ordinal preferences. Given a mixed strategy profile $x$, let us associate a vector $p_{w}(x)\in\mathbb{R}_{+}^{|M|}$ to each woman $w$ whose coordinates are sorted in the order of $w$’s preferences, such that if $m_{1}>_{w}m_{2}>_{w}\ldots>_{w}m_{n}$, we let the $j$th coordinate of $p_{w}(x)$ be given by
$p_{jw}(x)=x_{m_{j}w}\prod_{k>_{w}m_{j}}(1-x_{kw}),$
which is the probability that woman $w$ is matched to her $j$th favorite man, assuming that the players propose to women independently and according to their mixed strategies $x$. We note that for a pure-strategy profile $x\in\{0,1\}^{|M||W|}$, $p_{w}(x)$ would be a binary vector with all entries being zero except the $j$th one being one, where $j$ is the index of the highest ranked man proposing to $w$. Therefore, Lemma 6 asserts that after each best response move, the vectors $\{p_{w}(x),w\in W\}$ increase lexicographically with at least one of the vectors increases strictly. Now, if we again encode woman $w$’s preferences in a cardinal vector $\lambda_{w}=(\lambda_{wm_{1}},\ldots,\lambda_{wm_{n}})$, then for a pure strategy profile $x\in\{0,1\}^{|M||W|}$, we have
@@ A-F An Alternative Form for the Potential Function in Lemma 6
which coincides with the definition of $r_{w}(x)$ in Theorem 6 that picks the reward of the highest-ranked man proposing to $w$ under $x$. Therefore, in the space of cardinal preferences and for any pure-strategy profile $x\in\{0,1\}^{|M||W|}$, one can rewrite the potential function in Theorem 6 in a compact as
@@
whose value increases after every best response update by a player. We note the similarity between the potential function (88) and sum of players’ payoffs $\sum_{m}u_{m}(x)=\sum_{m,w}\mu_{mw}x_{mw}\prod_{k>_{w}m}(1-x_{kw})$, where one is written using mens’ preferences $\{\mu_{mw}\}$ and the other using the womens’ preferences $\{\lambda_{wm}\}$.
A-G A Gap Estimation Oracle for Unknown $\Delta$.
@@
Algorithm 3 Gap Estimation Oracle for Player $m$

### v2

@@ VI Global Learning of Stable Matchings in General Markets
In a pure strategy profile $x=(x_{m},m\in M)$, a matched man is the one whose pure strategy $x_{m}$ has exactly one coordinate equal to 1 (say $w$th coordinate), and all other coordinates are zero, and no other man of higher preference proposes to $w$, i.e., $x_{kw}=0\ \forall k>_{w}m.$ A pure strategy profile $x$ is a good state if all the men matched under $x$ are at their best responses. In other words, a good state $x$ refers to a partial matching in which the matched men are at their best responses.
@@
According to the payoff structures, the best response of player $m$ at time $t$ is to propose to his most preferred achievable woman $w^{\prime}$, i.e., among all women who rank $m$ higher or equal to their match under $x^{t-1}$, $m$ proposes to the woman $w^{\prime}$ which brings him the highest utility $\mu_{mw^{\prime}}$. Note that in particular $\mu_{mw^{\prime}}>\mu_{mw}$. Thus, the first time $\tau\leq t-1$ when man $m$ proposed and got matched to woman $w$, the woman $w^{\prime}$ was his more preferred choice and the reason why $m$ did not propose to $w^{\prime}$ at time $\tau$ was because $w^{\prime}$ was not achievable, i.e., $x^{\tau-1}_{m^{\prime}w^{\prime}}=1$ for some man $m^{\prime}>_{w^{\prime}}m$. Since $w^{\prime}$ was not achievable for $m$ in $x^{\tau-1}$ but she is achievable in $x^{t-1}$ implies that $r_{w^{\prime}}(x^{t-1})\!<\!r_{w^{\prime}}(x^{\tau-1})$. Therefore, $r_{w^{\prime}}(x^{\tau^{\prime}})\!<\!r_{w^{\prime}}(x^{\tau^{\prime}-1})$ for some $\tau\leq\tau^{\prime}\leq t-1$, contradicting the choice of $t$Q.E.D.
In fact, one can view the DA algorithm of [1] as a special version of the above lemma in which all the players are selected to best respond at each round, and the initial good state is the empty matching. Lemma 6 shows that the stable matching game admits an ordinal potential function if the game starts from a good state (we refer to Appendix B-D for another interpretation of this potential function). However, that does not imply that the stable matching game is an ordinal potential game [21], which requires the existence of a potential function whose changes are aligned with players’ payoff changes from any initial state. Nonetheless, the existence of such a “weak” potential function allows us to show the weak acyclicity of the stable matching game as defined next.
@@
In a weakly acyclic game, an arbitrarily high probability convergence can be guaranteed to a strict NE under appropriate learning dynamics [23]. If we do not have strictness property, the learning dynamics may oscillate between two nonstrict NE points with a positive probability. Using the above definitions, we are now ready to state the main result of this section, which provides a decentralized and uncoordinated algorithm for learning stable matchings in general markets with unknown preferences. The algorithm is an adaptation of the sample experimentation algorithm for weakly acyclic games [23] to the stable matching game. Intuitively, each player $m$ goes through a sequence of phase-dependent exploration/exploitation. At the end of each episode, player $m$ evaluates his payoff for each action he has taken during the last episode. He then updates his baseline action (an action that brings him the highest reward in the past episode up to some tolerance) and sticks to that baseline action most of the time in the next episode while still exploring new actions with a small probability. The detailed description of the algorithm is summarized in Algorithm 2.
@@
  • During the $s$th episode, player $m$ selects his baseline action obtained at the end of the previous episode with probability $1-\epsilon$ or explores a new uniformly sampled action with probability $\epsilon$:
@@
  • At the end of episode $s$, player $m$ evaluates his average utility when playing each action as
@@
Next, we note that any pure NE in the stable matching game is a strict NE. This follows directly by the fact that there are no ties between men’s preferences in the stable matching game. More precisely, let $\alpha^{*}$ be a pure NE (stable matching) and assume $u_{m}(\alpha^{*})=\mu_{mw}$. For any $\alpha_{m}\neq\alpha^{*}_{m}$, either $u_{m}(\alpha_{m},\alpha^{*}_{-m})$ equals 00 or $\mu_{mw^{\prime}}$ for some $w^{\prime}\neq w$. Because $\alpha^{*}_{m}$ is the best response of player $m$ at the NE and there are no ties between man $m$’s preferences, we must have $\mu_{mw^{\prime}}<\mu_{mw}$. Since $0<\mu_{mw}$, in either case we have $u_{m}(\alpha_{m},\alpha^{*}_{-m})<u_{m}(\alpha^{*}_{m},\alpha^{*}_{-m})$.
@@
VII Conclusions
In this work, we considered the problem of learning stable matchings with unknown preferences in a fully decentralized and uncoordinated manner. Using a game-theoretic formulation, we showed how to leverage the rich literature for learning NE in noncooperative games to design principled learning algorithms for driving the matching market to its stable points. In particular, we established several global and local convergence results with performance guarantees for learning stable matchings in hierarchical markets, general markets, and under weak/strong information feedback. Our results provide new insights for learning stable matchings through NE learning in noncooperative games and bridge the discrete problem of learning a stable matching with learning NE in continuous-action games.
This work opens several avenues for future research, some of which are discussed below.
  • We showed in Theorem 3 that for hierarchical markets it is possible to learn a stable matching with logarithmic regret using a fully decentralized and uncoordinated algorithm. However, the regret bound includes a constant factor that scales exponentially with the size of the market $n$. While we believe this exponential factor is necessary when using the EXP algorithm, an interesting question is whether there exist other fully decentralized and uncoordinated algorithms that can eliminate this exponential factor from the regret bound. More generally, it is important to characterize the information-theoretic limits above and below which efficient learning of a stable matching in a fully decentralized and uncoordinated fashion may or may not be possible.
  • In Theorem 5, we showed that if the EXP algorithm converges with positive probability, it must converge to a stable matching. We conjecture that, for general markets, Algorithm 1 converges globally to a stable matching with arbitrarily high probability. In that case, Algorithm 1 provides a simpler decentralized and uncoordinated alternative to Algorithm 2 that globally converges to a stable matching in general matching markets.
  • The waiting list feedback provided in Appendix A is merely an initial attempt to design faster decentralized and uncoordinated algorithms under richer information feedback. Extending our results to include other types of information feedback and studying their effect on the convergence speed of the market to its stable points would be another interesting research direction.
Appendix A Matching Markets with Stronger Information Feedback
So far in our work, we have assumed that men receive the minimum amount of information feedback as they interact with the market; namely, men only get to observe whether their proposal gets rejected or, if matched, they observe a random realization of their preference. In this appendix, we consider the learning problem from a new perspective by addressing the following question: Is there stronger information feedback that will allow us to derive the market faster globally to a stable matching in a decentralized and uncoordinated fashion? Therefore, we address the problem of learning a stable matching from an information feedback design perspective by showing that if the players are allowed to receive a bit more information as feedback, then the stable matching game will become a monotone game, hence allowing a wide range of decentralized and uncoordinated learning algorithms for obtaining its NE points.
@@ Appendix A Matching Markets with Stronger Information Feedback
Consider the regularized stable matching game with list feedback, where each player aims to maximize its regularized payoff function given by
$U_{m}(x_{m},x_{-m}):=u_{m}(x_{m},x_{-m})-\frac{\beta}{2}\|x_{m}\|^{2}.$
Then, for any two strategy profiles $x^{\prime},x$, we have
@@
The monotonicity property in Theorem 8 holds when the regularizer parameter $\beta$ is sufficiently large. The regularizer implies that players tend to maximize their approximate payoff functions. Thus, instead of learning a NE for the payoff functions (81), they learn an $\epsilon$-NE where the approximation error $\epsilon$ grows as $\beta$ increases. As a result, applying decentralized learning algorithms to learn a NE of the regularized game only guarantees an $\epsilon$-stable matching in the sense that the players will achieve a solution in which each player would be $\epsilon$ worse off than what he could achieve in a stable matching. Therefore, the regularizer parameter trades off between the market’s convergence speed and the accuracy of the learned stable matching. Finally, we note that the assumption of waiting list feedback is not necessary for faster convergence in any general market; it is merely one choice that we consider in this work, and there could be other types of information feedback that allow faster convergence. Nevertheless, our game-theoretic approach to the stable matching problem provides valuable insights on how to properly design the utilities and the information feedback to devise faster decentralized and uncoordinated algorithms for learning stable matchings.
Appendix B Supplementary Materials and Omitted Proofs
B-A Freedman’s Concentration Inequality for Martingales
@@
The proof can be found in [24, Theorem 1.6]Q.E.D.
B-B Proof of Lemma 4
@@ B-C Proof of Lemma 5
B-D An Alternative Form for the Potential Function in Lemma 6
One can represent the potential function in Lemma 6 directly for womens’ ordinal preferences. Given a mixed strategy profile $x$, let us associate a vector $p_{w}(x)\in\mathbb{R}_{+}^{|M|}$ to each woman $w$ whose coordinates are sorted in the order of $w$’s preferences, such that if $m_{1}>_{w}m_{2}>_{w}\ldots>_{w}m_{n}$, we let the $j$th coordinate of $p_{w}(x)$ be given by $p_{jw}(x)=x_{m_{j}w}\prod_{k>_{w}m_{j}}(1-x_{kw}),$ which is the probability that woman $w$ is matched to her $j$th favorite man, assuming that the players propose to women independently and according to their mixed strategies $x$. We note that for a pure-strategy profile $x\in\{0,1\}^{|M||W|}$, $p_{w}(x)$ would be a binary vector with all entries being zero except the $j$th one being one, where $j$ is the index of the highest ranked man proposing to $w$. Therefore, Lemma 6 asserts that after each best response move, the vectors $\{p_{w}(x),w\in W\}$ increase lexicographically with at least one of the vectors increases strictly. Now, if we again encode woman $w$’s preferences in a cardinal vector $\lambda_{w}=(\lambda_{wm_{1}},\ldots,\lambda_{wm_{n}})$, then for a pure strategy profile $x\in\{0,1\}^{|M||W|}$, we have
@@ B-D An Alternative Form for the Potential Function in Lemma 6
which coincides with the definition of $r_{w}(x)$ in Lemma 6 that picks the reward of the highest-ranked man proposing to $w$ under $x$. Therefore, in the space of cardinal preferences and for any pure-strategy profile $x\in\{0,1\}^{|M||W|}$, one can rewrite the potential function in Theorem 6 in a compact as
@@
whose value increases after every best response update by a player. We note the similarity between the potential function (100) and sum of players’ payoffs $\sum_{m}u_{m}(x)=\sum_{m,w}\mu_{mw}x_{mw}\prod_{k>_{w}m}(1-x_{kw})$, where one is written using mens’ preferences $\{\mu_{mw}\}$ and the other using womens’ preferences $\{\lambda_{wm}\}$.
B-E A Gap Estimation Oracle for Unknown $\Delta$.
@@
Algorithm 3 A Gap Estimation Oracle for Player $m$